OpenStudy (anonymous):

can someone help me please find an equation of the line. write the equation using function notation. through (4,-9); perpendicular to 3y=x-6

OpenStudy (anonymous):

3y=x-6 \[y=\frac{ 1 }{ 3 }x-2\] its slope=1/3 slope of line perpendicular to it=-3 reqd line is\[y-(-9)=-3(x-4)\] simplify it.

OpenStudy (anonymous):

so, y+9=-3x+12?

OpenStudy (anonymous):

y+-3x+3?

OpenStudy (anonymous):

y=-3x+3

OpenStudy (anonymous):

correct